The dimensions of a/b in the equation P = (a-t^2)/bx where P is pressure, x is distance and t is time are
    (a)    M2LT–3
    (b)    MT–2
    (c)    ML3T–1
    (d)    LT–3

Grade:upto college level

4 Answers

Simran Bhatia
10 years ago

(b)

P =  Þ Pbx = a – t2 ® [Pbx] = [a] = [T2]
Or [b] =  = [M–1T4]
= [MT–2]
Swetank Gupta
7 years ago
[a]=[t2]=T
[P]=T2/[bx]
[b]=T2/[P][x]=T2/ML-1T-2L=M-1T4
a/b=T2/M-1T4=MT-2
Thanks And Regards
Swetank Gupta
Karan Kaushal
6 years ago
The ratio of power out put of engines of two helicopters when linear size of one helicopters is one fourth of linear size of other and all other parameters are same is.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

[a]=[t2]=T2
[P]=T2/[bx]
[b]=T2/[P][x]=T2/ML-1T-2L=M-1T4
a/b=T2/M-1T4=MT-2

Thanks and Regards